OpenStudy (priyar):

A uniform wire (Y-given) is subjected to longitudinal tensile stress of 'S'. the overall volume change in the wire is 0.02% the fractional decrease in the radius of the wire is?

OpenStudy (priyar):

@ganeshie8 @samigupta8 @Rushwr

OpenStudy (priyar):

i am getting an answer..but its not correct..

OpenStudy (priyar):

@ParthKohli

Parth (parthkohli):

Hi :)

OpenStudy (priyar):

ok..i did like this: (though its not correct! pls tell me where i went wrong..) first i used stress/strain = Y i found (delta L/ L) and then, \[(\frac{ \Delta r }{ r })^2 = (\Delta V / V)*(L/\Delta L)\] then i got the req quantity from this..

OpenStudy (priyar):

its cylindrical..not spherical..right?

OpenStudy (priyar):

yes..

OpenStudy (priyar):

The values of Y ans S are given..

OpenStudy (priyar):

no..

OpenStudy (priyar):

this is the full Q except for the numerical values..

Parth (parthkohli):

\[Y\rm \Delta l/l = S\]\[\rm \Rightarrow l' = l\left(1 + \frac{S}Y\right)\]\[\rm \Delta V = \pi r^2 l - \pi r'^2 l' \]\[\Rightarrow \rm \frac{\Delta V}{V}= 1 - \frac{r'^2 l'}{r^2 l }\]\[\Rightarrow \rm \frac{\Delta V}{V} = 1- \frac{(r - \Delta r)^2(1 + S/Y)}{r^2}\]\[\Rightarrow \rm \frac{\Delta V}V = 1 - (1 + S/Y)\left(1 - \frac{\Delta r}r\right)^2\]You're given \(\rm \Delta V/V\) and you require \(\rm -\Delta r/r\). Cheers!

Parth (parthkohli):

Yes, I've considered everything in that answer.

OpenStudy (irishboy123):

wondering if you might also go down this road.... \(V = \pi r^2 l\) \(\delta V = 2 \pi r l \; \delta r + \pi r^2 \delta l\) \(\dfrac{\delta V }{ V} = \dfrac{2 \pi r l \; \delta r}{\pi r^2 l} + \dfrac{\pi r^2 \delta l}{\pi r^2 l} = 2 \dfrac{\delta r }{r} + \dfrac{S}{Y} \) all we've done here is linearise

OpenStudy (priyar):

thats a very nice idea! i got the answer! but i have a doubt why overall volume change means delta V/V and not just delta V?